# Neuronal Attention Circuit (NAC) for Representation Learning

### v2

@@ 2.3 NAC as Universal Approximator
We now establish the universal approximation capability of NAC by extending the classical Universal Approximation Theorem (UAT) (nishijima2021universal) to the proposed mechanism. For brevity, we consider a network with a single NAC layer processing fixed-dimensional inputs, though the argument generalizes to sequences.
@@
Let $K\subset\mathbb{R}^{n}$ be a compact set and $f:K\to\mathbb{R}^{m}$ be a continuous function. For any $\epsilon>0$, there exists a neural network consisting of a single NAC layer, with sufficiently large model dimension $d_{\text{model}}$, number of heads $H$, sparsity $s$, and nonlinear activations, such that the network’s output $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfies
@@ 3 Evaluation
We evaluate the proposed architecture against a range of baselines, including (DT & CT) RNN, (DT & CT) attention, and multiple NAC ablation configurations. Experiments are conducted across diverse domains, including irregular time-series modeling, lane keeping of autonomous vehicles, and Industry 4.0 prognostics. All results are obtained using 5-fold cross-validation, where models are trained using BPTT (see Appendix C.2) on each fold and evaluated across all folds. We report the mean ($\mu$) and standard deviation ($\sigma$) to capture variability and quantify uncertainty in the predictions. Table 1 provides results for all experiments, and the details of the baselines, ablation, environment utilized, the data curation and preprocessing, and neural network architectures for all experiments are provided in the Appendix D.3.
@@
Model Irregular Time-Series Lane-Keeeping of AVs Industry 4.0
E-MNIST (↑)
PAR (↑) CarRacing (↑)
Udacity (↓)
PRONOSTIA (↓) XJTU-SY (↓) HUST (↓)
@@ 3.1 Irregular Time-series
Event-based MNIST: Event-based MNIST is the transformation of the widely recognized MNIST dataset with irregular sampling added originally proposed in (lechner2022mixed). The transformation was done in two steps: (i) flattening each 28×28 image into a time series of length 784, and (ii) encoding the binary time series into an event-based format by collapsing consecutive identical values (e.g., 1,1,1,1 → (1, t=4)). This representation requires models to handle temporal dependencies effectively. NAC-PW achieved first place with an accuracy of 96.64%, followed by NAC-Exact/05s/8k at 96.12%. GRU-ODE and ContiFormer ranked third with 96.04%.
Person Activity Recognition (PAR): We employed the Localized Person Activity dataset from UC Irvine (localization_data_for_person_activity_196). The dataset contains data from five participants, each equipped with inertial measurement sensors sampled every 211 ms. The goal of this experiment is to predict a person’s activity from a set of predefined actions, making it a classification task. All models performed well on this task, with NAC-PW achieving 89.15% accuracy and taking first place. NAC-Exact/05s/8k and GRU-ODE ranked second with 89.01% accuracy, while NAC-02s ranked third with 88.84% mean accuracy.
3.2 Lane-Keeping of Autonomous Vehicles
Lane keeping in autonomous vehicles (AVs) is a fundamental problem in robotics and AI. Early works (tiang2018lane; park2021convolutional) primarily emphasized accuracy, often relying on large models. More recent research (lechner2020neural; razzaq2023neural) has shifted toward designing compact architectures suitable for resource-constrained devices. The goal of this experiment is to create a long causal structure between the road’s horizon and the corresponding steering commands. To evaluate, we used two widely adopted simulation environments: (i) OpenAI CarRacing (brockman2016openai); and (ii) the Udacity Self-Driving Car Simulator (udacity). In OpenAI CarRacing, the task is to classify steering actions from a predefined action set. In contrast, the Udacity Simulator requires predicting a continuous trajectory of steering values. We implemented the AI models proposed by (razzaq2023neural), replacing the recurrent layer with NAC and its counterparts. All models achieved around 80% accuracy on average in the CarRacing benchmark. Notably, NAC-PW performed the best, reaching the highest accuracy of 80.72%, followed by NAC-Steady, ranked second with 80.62%. LSTM and GRU took third position, achieving 80.60% on average. In the Udacity benchmark, NAC-32k performed the best, achieving the lowest MSE of 0.0170. NAC-Exact followed with 0.0173, and ContiFormer ranked third with 0.0174. To visualize saliency maps for these experiments, refer to Appendix D.3.3. Experimental videos are available for the OpenAI CarRacing [click here] and for the Udacity Simulator [click here].
@@ 3.2 Lane-Keeping of Autonomous Vehicles
Industry 4.0 has transformed manufacturing, making prognostic health management (PHM) systems essential. A key PHM task is estimating the remaining useful life (RUL) of components, particularly rolling element bearings (REB), which account for 40–50% of machine failures (ding2021remaining; zhuang2021temporal). The objective is to learn degradation features from one operating condition of a dataset and generalize to unseen conditions within the same dataset. Furthermore, the model should provide accurate RUL estimation on entirely different datasets, while maintaining a compact architecture suitable for resource-constrained devices, thereby supporting localized safety.
@@
We utilized three benchmark datasets: (i) PRONOSTIA (nectoux2012pronostia), (ii) XJTU-SY (wang2018xjtu), and (iii) HUST (thuan2023hust). Training is performed on PRONOSTIA, while XJTU-SY and HUST are used to assess cross-validation. We used the Score metric (nectoux2012pronostia) to assess the performance. We evaluate generalization using Bearing 1 from the first operating condition of each dataset. Figure 3 visualizes the expected degradation alongside the outputs of NAC. On the PRONOSTIA dataset, ContiFormer performed the best with the lowest score of 27.82. NAC-Exact/05s/8k and NAC-PW achieved nearly identical scores, obtaining 37.75 and 37.50 on average, respectively. On the XJTU-SY dataset, NAC-Exact/05s/8k has the lowest score of 19.87. NAC-FC ranked second with a score of 24.89, followed by NAC-PW in third place with an average score of 28.01. A similar trend was observed on the HUST dataset, where NAC-Exact/05s/8k achieved first place with a score of 27.82, NAC-PW ranked second with 30.14, and NAC-FC ranked third with 35.35. These results demonstrated the strong cross-validation capability of NAC.
Table 2: Run-Time and Memory Benchmark Results
Model
Run-Time
Throughput
(seq/s)
Peak Memory
RNN $1.8392^{\scriptsize\pm 0.1933}$ 0.544 0.29
CT-RNN $7.1097^{\scriptsize\pm 0.3048}$ 0.141 0.67
LSTM $2.6241^{\scriptsize\pm 0.2906}$ 0.381 0.42
PhasedLSTM $4.9812^{\scriptsize\pm 0.272}$ 0.201 0.80
GRU $3.216^{\scriptsize\pm 0.2566}$ 0.311 0.54
GRU-ODE $12.2498^{\scriptsize\pm 0.0525}$ 0.082 0.64
mmRNN $7.5852^{\scriptsize\pm 0.2785}$ 0.132 0.96
LTC $14.643^{\scriptsize\pm 0.2445}$ 0.068 0.99
CfC $6.0988^{\scriptsize\pm 0.2135}$ 0.164 0.76
Attention $0.0016^{\scriptsize\pm 0.0001}$ 625.00 16.86
MHA $0.0041^{\scriptsize\pm 0.0001}$ 243.90 69.05
mTAN $0.0272^{\scriptsize\pm 0.0054}$ 36.76 790.16
ODEFormer $0.0317^{\scriptsize\pm 0.0016}$ 31.55 67.71
CTA $8.5275^{\scriptsize\pm 0.2355}$ 0.117 1.43
ContiFormer $0.066^{\scriptsize\pm 0.0075}$ 15.15 67.71
NAC-2k $7.3071^{\scriptsize\pm 0.1547}$ 0.137 44.75
NAC-32k $7.2313^{\scriptsize\pm 0.219}$ 0.138 549.86
NAC-PW $8.5649^{\scriptsize\pm 0.0203}$ 0.117 5042.09
NAC-FC $0.0195^{\scriptsize\pm 0.0002}$ 51.28 29.92
NAC-02s $7.252^{\scriptsize\pm 0.2018}$ 0.138 151.54
NAC-09s $7.222^{\scriptsize\pm 0.176}$ 0.139 150.85
NAC-Exact/05s/8k $7.4101^{\scriptsize\pm 0.1586}$ 0.135 151.50
NAC-Euler $7.3367^{\scriptsize\pm 0.1719}$ 0.136 152.22
NAC-Steady $7.2942^{\scriptsize\pm 0.1451}$ 0.137 150.86
3.4 Run Time and Memory Experiments
We evaluate computational requirements on fixed-length sequences of 1024 steps, 64-dimensional features, 4 heads, and a Batch size of 1. Each model is run for ten forward passes on Google Colab T4-GPU, and we report the mean runtime with standard deviation, throughput, and peak memory usage. NAC occupies an intermediate position in runtime relative to several CT-RNN models, including GRU-ODE, CfC, and LTC. In terms of memory consumption, NAC uses significantly less memory than mTAN, with NAC-2k being the least memory-consuming among the CT-Attention models. Reducing NAC sparsity from 90% to 20% has minimal effect on memory, decreasing usage slightly from 151.54 MB to 150.85 MB. In constrast, decreasing the Top-K selection from PW to $k=2$ drastically reduces memory consumption from 5042 MB to 44.75 MB, demonstrating the flexibility of NAC.
@@ 3.4 Run Time and Memory Experiments
Interpreting the Results: We observe that increasing sparsity improves system robustness and leads to higher overall accuracy. Similarly, increasing Top-$K$ interactions improves accuracy, though the benefits diminish with growing memory consumption. We recommend a Top-$K$ of 8 with a 50% threshold as a robust default for this trade-off. In general, for tasks requiring adaptive continuous-time dynamics due to evolving temporal patterns, use the transient mode (Exact or Euler). For tasks where temporal dynamics are negligible, use the Steady mode for fast inference.
@@ 4 Discussions
Architectural improvement: Currently, NAC uses pre-determined wiring (AutoNCP) requiring three inputs: number of units (sensory + interneuron + motor), output motor neurons, and sparsity, with typically 60% of units assigned to sensory neurons. To integrate with the attention mechanism while preserving wiring, sensory units for $\mathcal{NN_{\text{sensory}}}$ are set as $\text{units}_{\text{sensory}}=\left\lceil\frac{d_{\text{model}}-0.5}{0.6}\right\rceil$ and backbone units as $\text{units}_{\textit{backbone}}=d_{\text{model}}+\left\lfloor\frac{d_{\text{model}}}{0.6}\right\rfloor$, where $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ denote the ceiling and floor functions, respectively. This results in a larger overall architectural size and increased runtime. Future work will support user-defined NCPs configurations or randomized wiring to enable more efficient architectures.
Learnable sparse Top-K selection: Sparse Top-K attention can miss important context, is sensitive to $k$, and may be harder to optimize. A further limitation is that it still computes the full $QK^{\top}$ matrix, which can dominate the cost for very long sequences. Future work includes adaptive or learnable Top-K selection, improved key scoring, and hardware-aware optimization to strengthen accuracy and robustness.
@@ 5 Conclusion
In this paper, we introduce the Neuronal Attention Circuit (NAC), a biologically inspired attention mechanism that reformulates attention logits as the solution to a first-order ODE modulated by nonlinear, interlinked gates derived from repurposing C.elegans nematode NCPs. NAC bridges discrete attention with continuous-time dynamics, enabling adaptive temporal processing without the limitations inherent to traditional scaled dot-product attention. Based on the solution to ODE, we introduce three computational modes: (i) Euler, based on explicit Euler integration; (ii) Exact, providing closed-form solutions; and (iii) Steady, approximating equilibrium states. In addition, a sparse Top-K pairwise concatenation scheme is introduced to mitigate the memory intensity. Theoretically, we establish NAC’s log-state stability, exponential error bounds, and universal approximation, thereby providing rigorous guarantees of convergence and expressiveness. Empirical evaluations demonstrate that NAC achieves state-of-the-art performance across diverse tasks, including irregularly sampled time-series benchmarks, autonomous vehicle lane-keeping, and industrial prognostics. Moreover, NAC occupies an intermediate position between CT-RNNs and CT-Attention, offering robust temporal modeling while requiring less runtime than CT-RNNs and less memory than CT-Attention models.
@@
The code for reproducibility is available at https://github.com/itxwaleedrazzaq/neuronal_attention_circuit
@@
The work addresses the growing field of continuous-time attention and pioneers a biologically plausible mechanism. It encourages research into sparse, adaptive networks that resemble natural wiring. From a societal perspective, it supports more robust AI in resource-limited settings, but it also raises ethical concerns when applied to areas such as surveillance or autonomous systems.
@@ Appendix A Preliminaries
A.1 Attention Mechanism
Attention mechanisms have become a cornerstone in modern neural architectures, enabling models to dynamically focus on relevant parts of the input. The concept was first introduced in the context of neural machine translation, where it allowed the decoder to weight encoder outputs according to their importance for generating each target token. Formally, given a query vector $q\in\mathbb{R}^{d}$, key vectors $K=[k_{1},k_{2},\dots,k_{n}]\in\mathbb{R}^{n\times d}$, and value vectors $V=[v_{1},v_{2},\dots,v_{n}]\in\mathbb{R}^{n\times d}$, the attention mechanism can be expressed in two steps:

### v3

@@ 2.2 Designing the Neural Network
2.2.1 Extension to Multihead
To scale this mechanism to multihead setting, we project the input sequence into $H$ independent subspaces (heads) of dimension $d_{\text{model}}/H$, yielding query, key, and value tensors $(q^{(h)},k^{(h)},v^{(h)})$ for $h\in\{1,\dots,H\}$. For each head, pairwise logits are computed according to Eqns. 2,3 or 18, followed by Softmax normalization to calculate attention weights. The resulting attention weights $\alpha^{(h)}_{t}$ are then integrated with the value vector $v^{(h)}$, to produce head-specific attention outputs. Finally, these outputs are concatenated and linearly projected back into the model dimension. This formulation ensures that each head learns distinct dynamic compatibilities governed by its own parameterization of $\phi$ and $\omega_{\tau}$, aggregation across heads preserves the expressive capacity of the standard multihead attention mechanism.
@@ 2.3 NAC as Universal Approximator
We now establish the universal approximation capability of NAC by extending the classical Universal Approximation Theorem (UAT) (nishijima2021universal) to the proposed mechanism. For brevity, we consider a network with a single NAC layer processing fixed-dimensional inputs, although the argument generalizes to sequences.
@@
Let $K\subset\mathbb{R}^{n}$ be a compact set and $f:K\to\mathbb{R}^{m}$ be a continuous function. For any $\epsilon>0$, there exists a neural network consisting of a single NAC layer, with a sufficiently large model dimension $d_{\text{model}}$, number of heads $H$, sparsity $s$, and nonlinear activations, such that the network’s output $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfies
@@ 3 Evaluation
We evaluate the proposed architecture against a range of baselines, including (DT & CT) RNNs, (DT & CT) Attentions, and multiple NAC ablation configurations. Experiments are conducted across diverse domains, including irregular time-series modeling, the lane keeping of autonomous vehicles, and Industry 4.0 prognostics. All the results are obtained using 5-fold cross-validation, where the models are trained using BPTT (see Appendix C.3) on each fold and evaluated across all the folds. We report the mean ($\mu$) and standard deviation ($\sigma$) to capture variability and quantify uncertainty in the predictions. Table 1 provides the results for all the experiments, and the details of the baselines, ablations, environments, data, preprocessing, and neural network architectures for all the experiments are provided in Appendix D.3.
@@
Model Irregular Time-Series Lane-Keeping of AVs Industry 4.0
E-MNIST (↑) PAR (↑) CarRacing (↑) Udacity (↓) PRONOSTIA (↓) XJTU-SY (↓) HUST (↓)
[…]
@@ 3.1 Irregular time-series
Event-based MNIST: Event-based MNIST is the transformation of the widely recognized MNIST dataset with irregular sampling added originally proposed in (lechner2022mixed). The transformation was performed in two steps: (i) flattening each 28×28 image into a time-series of length 784 and (ii) encoding the binary time-series into an event-based format by collapsing consecutive identical values (e.g., 1,1,1,1 → (1, t=4)). This representation requires models to handle temporal dependencies effectively. NAC-PW achieves an accuracy of 96.64%, followed by NAC-Exact/05s/8k at 96.12%. GRU-ODE and ContiFormer ranked third with 96.04%.
Person Activity Recognition (PAR): We employed the Localized Person Activity dataset from UC Irvine (localization_data_for_person_activity_196). The dataset contains data from five participants, each equipped with inertial measurement sensors sampled every 211 ms. The goal of this experiment is to predict a person’s activity from a set of predefined actions, making it a classification task. All models performed well on this task, with NAC-PW achieving 90.18% accuracy and taking first place. NAC-Exact/05s/8k ranked second with 90.11% accuracy, while NAC-Steady ranked third with 89.79% mean accuracy.
3.2 Lane Keeping of Autonomous Vehicles
Lane keeping in autonomous vehicles (AVs) is a fundamental problem in robotics and AI. Early works (tiang2018lane) focused primarily on accuracy, often relying on large models. More recent research (lechner2020neural; razzaq2023neural) has shifted toward designing compact architectures suitable for resource-constrained devices. The goal of this experiment is to create a long structure between the road’s horizon and the corresponding steering commands. For evaluation, we used two widely adopted simulation environments: (i) OpenAI CarRacing (brockman2016openai); and (ii) the Udacity Self-Driving Car Simulator (udacity). In OpenAI CarRacing, the task is to classify steering actions from a predefined action set. In contrast, the Udacity Simulator requires the prediction of a continuous trajectory of steering values. We implemented the AI models proposed by (razzaq2023neural), replacing the recurrent layer with NAC and its counterparts. All models achieved approximately 80% accuracy on average in the OpenAI CarRacing benchmark. Notably, the NAC-PW performed the best, reaching the highest accuracy of 80.72%, followed by NAC-Steady, ranked second with 80.62%. LSTM and GRU take third position, achieving 80.60% on average. For the Udacity benchmark, NAC-32k performed the best, achieving the lowest MSE of 0.0170. NAC-Exact/05s/8k followed by 0.0173, and ContiFormer ranked third with 0.0174. Experimental videos are available for the OpenAI CarRacing [click here] and for the Udacity Simulator [click here].
@@ 3.2 Lane Keeping of Autonomous Vehicles
Industry 4.0 has revolutionized manufacturing, rendering prognostic health management (PHM) systems indispensable. A key PHM task is estimating the remaining useful life (RUL) of components, particularly rolling element bearings (REB), which account for 40–50% of machine failures (ding2021remaining; zhuang2021temporal). The objective is to learn degradation features from one operating condition of a dataset and generalize them to unseen conditions within the same dataset. Furthermore, the model should provide accurate RUL estimation on entirely different datasets while maintaining a compact architecture suitable for resource-constrained devices, thereby supporting localized safety.
@@
We utilized three benchmark datasets: (i) PRONOSTIA (nectoux2012pronostia), (ii) XJTU-SY (wang2018xjtu), and (iii) HUST (thuan2023hust). Training is performed on PRONOSTIA, while XJTU-SY and HUST are used to assess cross-validation. We used the Score metric (nectoux2012pronostia) to assess the performance. We evaluate generalization using Bearing 1 from the first operating condition of each dataset. Figure 3 visualizes the expected degradation alongside the outputs of NAC. For the PRONOSTIA dataset, NAC-Exact/05s/8k performed the best, with the lowest score of 37.75. NAC-PW and Contiformer achieved nearly identical scores, with average values of 37.75 and 37.82, respectively. For the XJTU-SY dataset, NAC-Exact/05s/8k has the lowest score of 22.87. CTA ranked second with 25.86, followed by NAC-$\phi_{\text{linear}}$ with 29.53. A similar trend was observed on the HUST dataset, where NAC-Exact/05s/8k achieved first place with a score of 27.82, NAC-PW ranked second with 30.14, and NAC-$\phi_{\text{linear}}$ ranked third with 35.89. These results demonstrated the strong cross-validation adaptability of NAC.
Figure 4: Intuitive visualization of NAC cell activity during driving. (a) NAC-controlled car on the driving map. (b) Sensory gate neurons output (centroid) for the $q$, $k$, and $v$ projections. (c) Backbone neuron activity (centroid), including interneurons, command neurons, and motor neurons. (d) $\omega_{\tau}$ and $\phi$ and $a_{t}$ computations. Internal plots show actual neuron activities.
3.4 NAC enhances Interpretability
Interpretability refers to the process of providing explanations that are understandable to humans, either through visualization or analysis at the neuron cell level. We analyze NAC’s interpretability on OpenAI CarRacing trained with NAC-Steady at two levels: (i) Saliency maps, which use Grad-CAM (selvaraju2017grad) to visualize which features are attended to by the convolutional layers of the model during inference. Additional information on saliency maps is provided in the Appendix D.3.3; (ii) Cell-level neuron state, where we analyze neuron outputs at the cell level, starting from the sensory gates, including the $q$, $k$, and $v$ projections, and extending through the backbone inter-command-motor neurons to the computation of the attention logits. Figure 4 illustrates the cell-level neuron states of NAC. We observe that $\omega_{\tau}$ couples its values between 0.8 and 1.6, depending on the input scene. When the car turns left or left-straight, the $\omega_{\tau}$ value remains below 1.1. When it turns right or right-straight, $\omega_{\tau}$ is maintained above 1.3. In intermediate cases, the model drives straight. This behavior indicates that the evolution of $\omega_{\tau}$ changes over time, confirming its input dependency. Similarly, the attention logits reflect this behavior, with a tighter range for left and left-straight actions (below 0.6), higher values for right and right-straight actions (above 0.65), and intermediate values corresponding to straight driving. These cell-level neuron states highlight how activations depend on the input, providing a way to understand and evaluate safety-critical decisions.
3.5 Interpreting the Results
@@
We observe that increasing sparsity improves system robustness and leads to higher overall accuracy. Similarly, increasing the number of Top-$K$ interactions improves accuracy, although the benefits diminish. We recommend a Top-$K$ of 8 with 50% sparsity as a robust default for this trade-off. In general, for tasks requiring adaptive CT dynamics due to evolving temporal patterns, use the transient mode (Exact or Euler). For tasks where temporal dynamics are negligible, use the Steady mode.
@@ 4 Discussions
Runtime and Memory Experiment: The runtime and memory experiments (see Appendix D.4) show that NAC occupies an intermediate position between CT-RNN and CT-Attention baselines in terms of runtime and memory consumption, which could be further improved.
Architectural improvement: Currently, NAC uses predetermined wiring (AutoNCP) requiring three inputs: the number of units (sensory + interneuron + motor), motor neurons, and sparsity, with typically 60% of the units assigned to sensory neurons. To integrate with the attention mechanism while preserving wiring, sensory units for $\mathcal{NN_{\text{sensory}}}$ are set as $\text{units}_{\text{sensory}}=\left\lceil\frac{d_{\text{model}}-0.5}{0.6}\right\rceil$ and backbone units as $\text{units}_{\textit{backbone}}=d_{\text{model}}+\left\lfloor\frac{d_{\text{model}}}{0.6}\right\rfloor$, where $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ denote the ceiling and floor functions, respectively. This results in a larger overall architectural size. Future work will support user-defined NCPs configurations or randomized wiring to enable more efficient architectures.
Learnable sparse Top-K selection: Sparse Top-K attention can miss important context, is sensitive to $k$, and may be harder to optimize. Future work will include learnable Top-K selection, improved key scoring, and hardware-aware optimization to increase accuracy and robustness.
@@ 5 Conclusion
In this paper, we introduce the Neuronal Attention Circuit (NAC), a biologically inspired attention mechanism that reformulates attention logits as the solution to a first-order ODE modulated by nonlinear, interlinked gates derived from repurposing C.elegans nematode NCPs. NAC bridges discrete attention with CT dynamics, enabling adaptive temporal processing without the limitations inherent to traditional scaled dot-product attention. Based on the solution to ODE, we introduce three computational modes: (i) Euler, based on explicit Euler integration; (ii) Exact, providing closed-form solutions; and (iii) Steady, approximating equilibrium states. In addition, an adaptable sparse Top-K pairwise concatenation mechanism is implemented to mitigate the quadratic memory footprint. Theoretically, we establish NAC’s logit-state stability and exponential and frozen approximation error bounds, thereby providing rigorous guarantees of convergence and expressiveness. Empirically, we demonstrated that NAC achieves state-of-the-art performance across diverse tasks, including irregularly sampled time-series benchmarks, autonomous vehicle lane-keeping, and industrial prognostics, while being interpretable and explainable. Moreover, NAC occupies an intermediate position between CT-RNNs and CT-Attentions in terms of runtime and memory consumption.
@@
The code for reproducibility is available at https://github.com/itxwaleedrazzaq/neuronal_attention_circuit.
@@
This work addresses the growing field of CT-Attention and pioneers a biologically inspired mechanism. It encourages research into sparse, adaptive networks that resemble natural wiring. From a societal perspective, it supports the development of more robust AI in resource-limited settings. Nevertheless, it also raises ethical concerns when applied to areas such as surveillance or autonomous systems.
@@ Appendix A Preliminaries
A.1 Neuronal Circuit Policies (NCPs)
A.2 Attention Mechanism
Attention mechanisms have become a cornerstone in modern neural architectures, enabling models to dynamically focus on relevant parts of the input. The concept was first introduced in the context of neural machine translation, where it allows the decoder to weight encoder outputs according to their importance for generating each target token. Given a query vector $q\in\mathbb{R}^{d}$, key vectors $K=[k_{1},k_{2},\dots,k_{n}]\in\mathbb{R}^{n\times d}$, and value vectors $V=[v_{1},v_{2},\dots,v_{n}]\in\mathbb{R}^{n\times d}$, the attention mechanism can be expressed in two steps: